# Gaussian Grouping: Segment and Edit Anything in 3D Scenes

### v1

@@ Abstract
The recent Gaussian Splatting achieves high-quality and real-time novel-view synthesis of the 3D scenes. However, it is solely concentrated on the appearance and geometry modeling, while lacking in fine-grained object-level scene understanding. To address this issue, we propose Gaussian Grouping, which extends Gaussian Splatting to jointly reconstruct and segment anything in open-world 3D scenes. We augment each Gaussian with a compact Identity Encoding, allowing the Gaussians to be grouped according to their object instance or stuff membership in the 3D scene. Instead of resorting to expensive 3D labels, we supervise the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by SAM, along with introduced 3D spatial consistency regularization. Comparing to the implicit NeRF representation, we show that the discrete and grouped 3D Gaussians can reconstruct, segment and edit anything in 3D with high visual quality, fine granularity and efficiency. Based on Gaussian Grouping, we further propose a local Gaussian Editing scheme, which shows efficacy in versatile scene editing applications, including 3D object removal, inpainting, colorization and scene recomposition. Our code and models will be at https://github.com/lkeab/gaussian-grouping.
Figure 1: Our Gaussian Grouping approach jointly reconstructs (column a) and segments (column b) anything in full open-world 3D scenes, with fine-grained instance and stuff level modeling. This enables versatile scene editing applications, such as 3D object removal (column c), 3D object inpainting (column d) and scene re-composition and object colorization (column e). Since the segmentation information is encapsulated in the 3D Gaussians, editing tasks such as 3D object removal, colorization and object location exchange can be performed directly w/o training, while inpainting only requires minutes of fine-tuning.
@@ 1 Introduction
Open-world 3D scene understanding is an essential challenge, with far-reaching implications for robotics, AR / VR, and autonomous driving. Given a set of posed RGB images, our goal is to learn an effective 3D representation that jointly reconstructs and segments anything in the 3D scene. The representation should easily support a wide range of downstream scene editing applications. For example, in Figure 1, the 3D object of the scene can be easily removed or inpainted, and the scene can be recomposed by exchanging multiple object locations.
While there has been remarkable progress in 2D scene understanding brought by SAM and its variants [15, 12, 58], their extension to 3D has been constrained. This is mostly by the availability and the labor-intensive process of creating 3D scene datasets. Existing methods [37, 8] rely on manually-labeled datasets, which are both costly and limited in scope, or requires accurately scanned 3D point clouds [42, 33] as input. This hinders the development of the 3D scene understanding approaches that can quickly generalize across various real-world scenes.
By taking multi-view captures, existing NeRFs-based methods [39, 25, 14, 17] lift 2D masks or distill CLIP [34] / DINO [3] features via neural fields rendering. However, due to the implicit and continuous representation of NeRF, these methods require expensive random sampling and are computationally intensive to optimize. Further, it is hard to directly adjust NeRF-based approaches for the downstream local editing tasks [17], because the learned neural networks, such as MLPs, cannot decompose each part or module in the 3D scene easily. Several methods [38, 4] combine NeRF or stable-diffusion [36] with the masks of SAM, but they focus on a single object without attaining more holistic scene understanding.
Alternative to NeRFs, the recently emerged 3D Gaussian Splatting [13] has shown impressive reconstruction quality with high training and rendering efficiency. It represents the 3D scene with an array of coloured and explicit 3D Gaussians, which are rendered into camera views for novel view synthesis. Nevertheless, Gaussian Splatting does not model object instance or semantic understanding. To achieve fine-grained scene understanding, we extend this approach beyond merely capturing the scene’s appearance and geometry to include the individual objects and elements that constitute the 3D environments.
@@
To further boost the grouping accuracy, besides the standard cross-entropy loss for identity classification, we also introduce an un-supervised 3D Regularization Loss by leveraging 3D spatial consistency. The loss enforces the Identity Encodings of the top $K$-nearest 3D Gaussians to be close in their feature distance. We find it helps for Gaussians inside the 3D object or heavily occluded to be supervised during rendering more sufficiently.
We highlight the advantages of Gaussian Grouping not only by providing high visual quality and fast training, but also through its downstream scene editing applications brought by our dense 3D segmentation. Since our model captures the 3D scenes as compositional structures, each Gaussian group within the scene operates independently, allowing for parts to be fully decoupled or separated. This decomposition is crucial in scenarios where individual components need to be identified, manipulated, or replaced without disrupting the entire scene structure. We also propose an efficient Local Gaussian Editing strategy based on the grouping. By the grouped 3D Gaussians, we show extensive 3D scene editing cases, such as in 3D scene re-composition, object inpainting or object removal, with both impressive visual effect and fine granularity.
@@ 2 Related Works
3D Gaussian Models 3D Gaussian Splatting [13] has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering. Several follow-up methods [26, 52, 53] extend it to dynamic 3D scenes by tracking of all dense scene elements [26] or deformation field modeling [52, 49]. Another stream of works [6, 43, 54] focus on the 3D content creation, which combine Gaussian Splatting with the diffusion-based models and show high-quality generation results. However, none of the existing Gaussian Splatting works enables object / stuff-level or semantic understanding of the 3D scene. Our Gaussian Grouping extends Gaussian Splatting from pure scene appearance and geometry modeling to also support open-wold and fine-grained scene understanding. We show that the grouped 3D Gaussians is an effective and flexible 3D representation in supporting a series of downstream scene editing applications.
@@
Radiance-based Open World Scene Understanding Existing semantic scene modeling approaches combined with radiance fields are NeRF-based [29]. Semantic-NeRF [60] initiated the incorporation of semantic information into NeRFs and facilitated the generation of semantic masks from novel views. Building upon it, subsequent research focusing on in-domain scene modeling, has expanded the scope by introducing instance modeling [20, 9, 47, 39] or encoding visual features [44, 17, 45, 46, 35] that support semantic delineation. Unlike our approach, most of these methods are designed for in-domain scene modeling and cannot generalize to open-world scenario.
@@
By distilling 2D features such as CLIP [34] or DINO [3], Distilled Feature Fields [17] explores embedding pixel-aligned feature vectors, and LERF [14] learns language field to help open-world 3D semantics. In contrast to our work, these open-world approaches are only limited in semantic segmentation (hard to separate similar objects of the same category) and cannot produce very accurate segmentation masks as shown in our experiment. To our knowledge, we propose the first Gaussian-based method to tackle open-world 3D scene understanding, where we show the advantages comparing to existing NeRF-based approaches [14, 17, 39] in segmentation quality, efficiency and good extension to downstream scene editing applications.
@@
SAM in 3D Segment Anything Model (SAM) [15] was released as a foundational vision model for zero-shot 2D segmentation. Several works lift SAM’s 2D masks to 3D segmentation via NeRF [4, 5] or 3D point cloud [51]. However, these NeRF-based approaches only focus on single / few objects of the 3D scene, while our Gaussian Grouping operates in automatic everything mode to attain the holistic understanding of each instance / stuff of the full scene.
@@
Radiance-based Scene Editing Manipulating 3D scenes via radiance field is challenging. For existing NeRF-based scene editing / manipulation works [57, 23, 11, 22], Clip-NeRF [48], Object-NeRF [50], and LaTeRF [30] design approaches to change and complete objects represented by NeRFs; however, their application scenario is limited to simple objects, rather than scenes with significant clutter and texture. Some other works specify bounding boxes [32, 56], to allow flexible compositing of various objects [59] or combining with physical simulation [21]. Mostly recently, 3D object inpainting is studied in SPIn-NeRF [31] and language-based scene editing is proposed in [10]. Compared to them, we apply our Gaussian Grouping to versatile scene editing tasks, and show benefits brought by the fine-grained scene modeling, where multiple scene editings can be easily superimposed on the same image. The design of Local Gaussian Editing makes the whole training / editing process both simple and efficient.
@@ 3 Method
Our work aims to build an expressive 3D scene representation, which not only models appearance and geometry, but also captures every instance and stuff identity of the scene. We design our method based on the recent 3D Gaussian Splatting [13], and extend it from pure 3D reconstruction to fine-grained scene understanding. Our approach, called Gaussian Grouping, is capable of: 1) modeling each 3D part of the scene with appearance, geometry together with their mask identities; 2) fully decomposing the 3D scene into discrete groups, e.g. representing different object instances; 3) allowing for fast training and rendering, while not diluting the original 3D reconstruction quality.
Gaussian Grouping effectively leverages the dense 2D mask proposals of SAM, and lifts them to segment anything in the 3D scene via radiance fields rendering. In Sec 3.1, we first briefly review the 3D Gaussian Splatting method on radiance field rendering. We then detailed the input data pre-processing steps in Section 3.2.1, and further described the propose Gaussian Grouping in Section 3.2.2. With the constructed 3D representation, finally, we show its advantages in downstream scene editing tasks by the efficient Local Gaussian Editing in Section 3.3.
Figure 2: The method pipeline of our Gaussian Grouping contains three main steps: (a) We first prepare the input by deploying SAM to automatically generate masks in everything mode for each view independently. (b) Then, to obtain the consistent masks IDs across training views, we take an universal temporal propagation model [7] to associate the mask labels and generate a coherent multi-view segmentation. (c) With the prepared training input, we jointly learn all properties of the 3D Gaussians, including their group Identity Encoding, by differentiable rendering. Our encoding is supervised by the 2D Identity Loss, leveraging the coherent segmentation views, and a 3D Regularization loss. We use color the denote object IDs across frames for input views. We omit the rendering process for other Gaussian parameters and the density control part for simplicity, as it is inherited from [13].
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting, as introduced by Kerbl et al. [13], encapsulates 3D scene information using a suite of 3D colored Gaussians. This technique has established its effectiveness in the reconstruction tasks, exhibiting high inference speeds and remarkable quality of reconstruction within timeframes on par with those of NeRF. Yet, its potential in scene understanding has not been thoroughly investigated. Our research reveals that 3D Gaussians hold considerable promise for the open-world and complex 3D scene understanding.
To represent the scene, each Gaussian’s property is characterized by a centroid $\mathbf{p}=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D size $\mathbf{s}\in\mathbb{R}^{3}$ in standard deviations, and a rotational quaternion $\mathbf{q}\in\mathbb{R}^{4}$. To allow fast $\alpha$-blending for rendering, we assign an opacity value $\alpha\in\mathbb{R}$ and a color vector $\mathbf{c}$ represented in the three degrees of spherical harmonics (SH) coefficients. These adjustable parameters are collectively symbolized by $S_{\Theta_{i}}$, where $S_{\Theta_{i}}=\{\mathbf{p}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}\}$ represents the set of parameters for the $i$-th Gaussian. To visualize and supervise these 3D Gaussians, Gaussian Splatting projects them onto the 2D image plane. It implements differentiable rendering and gradient-based optimization for each pixel in a sequence of Gaussians sorted in the depth order, for computing the final color and opacity.
@@
In this section, we describe the design of our Gaussian Grouping. To enable 3D Gaussians for fine-grained scene understanding, our key insight is that we preserve all attributes of the Gaussians (such as their number, color, opacity, and size) in its original setting, but add new Identity Encoding parameters (similar to the format of color modeling). This allows each Gaussian to be assigned to its represented instances or stuffs in the 3D scene.
3.2.1 Anything Mask Input and Consistency
To model open-world scenes, instead of relying on the expensive 3D labels for supervision, we train our approach by utilizing SAM’s 2D masks [15] as input. Concretely, we take a set of multi-view captures along with the automatically generated 2D segmentations by SAM, as well as the corresponding cameras calibrated via SfM [40]. SAM is adopted here because it shows strong zero-shot capabilities not limited by the training categories, and supports automatically segmenting everything in a wide-variety of images.
@@ 3.2.1 Anything Mask Input and Consistency
2D Image and Mask Input To prepare the input for Gaussian Grouping, in Figure 2 (a), we first deploy SAM to automatically generate masks for each image of the multi-view collection. The 2D masks are individually produced per image. Then, to assign each 2D mask a unique ID in the 3D scene, we need to associate the masks of the same identity across different views and obtain the total number $K$ of instances / stuffs the 3D scene.
@@
Identity Consistency across Views Instead of resorting to the cost-based linear assignment [39] during training, we treat the multi-view images of a 3D scene as a video sequence with gradually changing views. To achieve the 2D mask consistency across views, we employ a well-trained zero-shot tracker [7] to propagate and associate masks. This also provides the total number of mask identities in the 3D scene. We visualize associated 2D mask labels in Figure 2 (b). Comparing to the cost-based linear assignment proposed in [39], we find it simplifies the training difficulty while avoids repeatedly compute the matching relation in each rendering iteration, resulting over 60$\times$ speeding up. It also achieves better performance than the cost-based linear assignment, especially under the dense and overlapping masks by SAM.
3.2.2 3D Gaussian Rendering and Grouping
Conventional 3D Gaussian paramters for modeling the scene includes 3D center position (mean), 3D size (anisotropic covariance), opacity, and color (Spherical Harmonic coefficients). Separate operations are taken for the adaptive control on the Gaussian density. As in Sec 3.1, these properties are differentiable and can be easily projected to 2D splats for the whole scene optimization.
@@
Identity Encoding To generate 3D-consistent mask identities across views of the scene, we propose to group 3D Gaussians belong to the same instance / stuff. In addtion to the existing Gaussian properties, we introduce a new parameter, i.e., Identity Encoding, to each Gaussian. The Identity Encoding is a learnable and compact vector of length 16, which we find is sufficient to distinguish the different objects / parts in the scene with computation efficiency. During training, similar to Spherical Harmonic (SH) coefficients on representing color of each Gaussian, we optimize the introduced Identity Encoding vector to represent its instance ID of the scene in the same manner. However, different from the view-dependent appearance modeling of the scene, the instance ID is consistent across various rendering views. Thus, we set the SH degree of the Identity Encoding to 0, only to model its direct-current component. Unlike NeRF-based methods [39, 17, 14] designing extra semantic MLP layers, the Identity Encoding serves as the learnable property for each Gaussian to group the 3D scene.
@@
Grouping via Rendering To optimize the introduced Identity Encoding of each Gaussian, in Figure 2 (c), we render these encoded identity vectors into 2D images in a differentiable manner. We take the differentiable 3D Gaussian renderer from [13] and treat the rendering process similar to the color (SH coefficients) optimization in [13].
@@
3D Gaussian splatting adopts neural point-based $\alpha^{\prime}$-rendering [18, 19], where the influence weight $\alpha^{\prime}$ can evaluated in 2D for each pixel for each Gaussian. Following [13], the influence of all Gaussians on a single pixel location are computed by sorting the Gaussians in depth order and blending $\mathcal{N}$ ordered points overlapping the pixels [27, 29]:
@@
Grouping Loss After associating 2D instance labels across each training view (Sec 3.2.1), suppose there are $K$ masks in total in the 3D scene. To group each 3D Gaussian by the instance /stuff mask identities, we design the grouping loss $\mathcal{L}_{\text{id}}$ for updating the Identity Encoding of Gaussians with two components:
@@
2) 3D Regularization Loss: To further boost the grouping accuracy of Gaussians, besides the standard cross-entropy loss for indirect 2D supervision, we also introduce an unsupervised 3D Regularization Loss to directly regularize the learning of Identity Encoding $e_{i}$. 3D Regularization Loss leverages the 3D spatial consistency, which enforces the Identity Encodings of the top $k$-nearest 3D Gaussians to be close in their feature distance. This allows the 3D Gaussians inside the 3D object, or heavily occluded (not visible in nearly all training views) during the point-based rendering (Eq. 1) being supervised more sufficiently. In Eq. 3, we denote $F$ as softmax operation combined after linear layer $f$ (shared in computing the 2D Identity loss). We formalize the KL divergence loss with $m$ sampling points as,

### v2

@@ Abstract
The recent Gaussian Splatting achieves high-quality and real-time novel-view synthesis of the 3D scenes. However, it is solely concentrated on the appearance and geometry modeling, while lacking in fine-grained object-level scene understanding. To address this issue, we propose Gaussian Grouping, which extends Gaussian Splatting to jointly reconstruct and segment anything in open-world 3D scenes. We augment each Gaussian with a compact Identity Encoding, allowing the Gaussians to be grouped according to their object instance or stuff membership in the 3D scene. Instead of resorting to expensive 3D labels, we supervise the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by Segment Anything Model (SAM), along with introduced 3D spatial consistency regularization. Compared to the implicit NeRF representation, we show that the discrete and grouped 3D Gaussians can reconstruct, segment and edit anything in 3D with high visual quality, fine granularity and efficiency. Based on Gaussian Grouping, we further propose a local Gaussian Editing scheme, which shows efficacy in versatile scene editing applications, including 3D object removal, inpainting, colorization, style transfer and scene recomposition. Our code and models are at github.com/lkeab/gaussian-grouping.
Keywords:
3D Segment Anything Scene Editing Gaussian Splatting
footnotetext:  Project Lead
Figure 1: Our Gaussian Grouping jointly reconstructs (column a) and segments (column b) anything in full open-world 3D scenes, with fine-grained instance and stuff level modeling. This enables versatile scene editing applications, such as 3D object removal (column c), 3D object inpainting (column d, which first removes the 3D object and then inpaints the holes) and scene re-composition and object colorization (column e). Since the segmentation information is encapsulated in the 3D Gaussians, editing tasks such as 3D object removal, colorization and object location exchange can be performed directly w/o training, while inpainting only requires minutes of fine-tuning.
@@ 1 Introduction
Open-world 3D scene understanding is an essential challenge, with far-reaching implications for robotics, AR / VR, and autonomous driving. Given a set of posed RGB images, our goal is to learn an effective 3D representation that jointly reconstructs and segments anything in the 3D scene. The representation should easily support a wide range of downstream scene editing applications. For example, in Figure 1, the 3D object of the scene can be easily removed or inpainted, and the scene can be recomposed by exchanging object locations.
While there has been remarkable progress in 2D scene understanding brought by SAM and its variants [17, 14, 64], their extension to 3D has been constrained. This is mostly due to the availability and the labor-intensive process of creating 3D scene datasets. Existing methods [42, 8] rely on manually-labeled datasets, which are both costly and limited in scope, or require accurately scanned point clouds [47, 37] as input. This hinders the development of the 3D scene understanding methods that can quickly generalize across various real-world scenes.
By taking multi-view captures, existing NeRF-based methods [44, 28, 16, 19] lift 2D masks or distill CLIP [39] / DINO [3] features via neural fields rendering. However, due to the implicit and continuous representation of NeRF, these methods require expensive random sampling and are computationally intensive to optimize. Further, it is hard to directly adjust NeRF-based approaches for the downstream local editing tasks [19], because the learned neural networks, such as MLPs, cannot decompose each part or module in the 3D scene easily. Several methods [43, 4] combine NeRF or stable-diffusion [41] with the masks of SAM, but they only focus on a single object.
Alternative to NeRFs, the recently emerged 3D Gaussian Splatting [15] has shown impressive reconstruction quality with high training and rendering efficiency. It represents the 3D scene with an array of colored and explicit 3D Gaussians, which are rendered into camera views for novel view synthesis. Nevertheless, Gaussian Splatting does not model object instances or semantic understanding. To achieve fine-grained scene understanding, we extend this approach beyond merely capturing the scene’s appearance and geometry to include the individual objects and elements that constitute the 3D environments.
@@
To further boost the grouping accuracy, besides the standard cross-entropy loss for identity classification, we also introduce an un-supervised 3D Regularization Loss by leveraging 3D spatial consistency. The loss enforces the Identity Encodings of the top $K$-nearest 3D Gaussians to be close in their feature distance. We find it helps Gaussians inside the 3D object or heavily occluded to be supervised during training more sufficiently.
We highlight the advantages of Gaussian Grouping not only by providing high visual quality and fast training, but also through its downstream scene editing applications brought by our dense 3D segmentation. Since our model captures the 3D scenes as compositional structures, each group of 3D Gaussians operates independently, allowing for parts to be fully decoupled or separated. This decomposition is crucial in scenarios where individual components need to be identified, manipulated, or replaced w/o disrupting the entire scene structure.
Our contribution can be summarized as follows:
  1. 1.
    We introduce Gaussian Grouping, the first 3D Gaussian Splatting-based segmentation framework that lifts knowledge of SAM to 3D scene anything zero-shot segmentation without the need for 3D mask labels.
  2. 2.
    Gaussian Grouping supports various downstream tasks with our proposed Local Gaussian Editing scheme. Individual components are identified, manipulated, or replaced without disrupting the entire scene structure. By the grouped 3D Gaussians, we show extensive 3D scene editing cases, such as 3D scene re-composition, object inpainting, object removal and object style transfer, with both impressive visual effect and fine granularity.
  3. 3.
    Our training and rendering processes are swift, ensuring compliance with real-time operational requirements.
@@ 2 Related Works
3D Gaussian Splatting [15] has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering. Several follow-up methods [29, 59, 58] extend it to dynamic 3D scenes by tracking of all dense scene elements [29] or deformation field modeling [59, 54]. Another stream of works [6, 48, 60] focuses on the 3D content creation, which combines Gaussian Splatting with diffusion-based models and shows high-quality generation results. However, none of the existing Gaussian Splatting works enables object / stuff-level or semantic understanding of the 3D scene. Our Gaussian Grouping extends Gaussian Splatting from pure scene appearance and geometry modeling to also support open-world and fine-grained scene understanding. We show that the grouped 3D Gaussians is an effective and flexible 3D representation in supporting a series of downstream scene editing applications.
2.0.2 Radiance-based Open World Scene Understanding
@@
Existing semantic scene modeling approaches combined with radiance fields are NeRF-based [33, 31, 66]. Semantic-NeRF [67] initiated the incorporation of semantic information into NeRFs and facilitated the generation of semantic masks from novel views. Building upon it, subsequent research focusing on in-domain scene modeling, has expanded the scope by introducing instance modeling [22, 11, 52, 44] or encoding visual features [49, 19, 50, 51, 40] that support semantic delineation. Unlike our approach, most of these methods are designed for in-domain scene modeling and cannot generalize to open-world scenarios.
@@
By distilling 2D features such as CLIP [39] or DINO [3], Distilled Feature Fields [19] explores embedding pixel-aligned feature vectors, and LERF [16] learns language field to help open-world 3D semantics. In contrast to our work, these open-world approaches are only limited in semantic segmentation (hard to separate similar objects of the same category) and cannot produce very accurate segmentation masks as shown in our experiment. To our knowledge, we propose the first Gaussian-based method to tackle open-world 3D scene understanding, where we show the advantages compared to existing NeRF-based approaches [16, 19, 44] in segmentation quality, efficiency and good extension to downstream scene editing applications.
2.0.3 SAM in 3D
@@
Segment Anything Model (SAM) [17] was released as a foundational vision model for zero-shot 2D segmentation. Several works lift SAM’s 2D masks to 3D segmentation via NeRF [4, 5] or 3D point cloud [57]. However, these NeRF-based approaches only focus on a single / few objects of the 3D scene, while our Gaussian Grouping operates in automatic everything mode to attain the holistic understanding of each instance / stuff of the full scene.
2.0.4 Radiance-based Scene Editing
@@
Manipulating 3D scenes via a radiance field is challenging. For existing NeRF-based scene editing / manipulation works [63, 27, 55, 13, 25], Clip-NeRF [53], Object-NeRF [56], and LaTeRF [35] design approaches to change and complete objects represented by NeRFs; however, their application scenario is limited to simple objects, rather than scenes with significant clutter and texture. Some other works specify bounding boxes [36, 62], to allow flexible compositing of various objects [65] or combining with physical simulation [24]. Most recently, 3D object inpainting is studied in SPIn-NeRF [34] and language-based scene editing is proposed in [12]. Compared to them, we apply our Gaussian Grouping to versatile scene editing tasks, and show benefits brought by the fine-grained scene modeling, where multiple scene editings can be easily superimposed on the same image. The design of Local Gaussian Editing makes the whole training / editing process both simple and efficient.
@@ 3 Method
Our work aims to build an expressive 3D scene representation, which not only models appearance and geometry, but also captures every instance and stuff identity of the scene. We design our method based on the recent 3D Gaussian Splatting [15], and extend it from pure 3D reconstruction to fine-grained scene understanding. Our approach, called Gaussian Grouping, is capable of: 1) modeling each 3D part of the scene with appearance, geometry together with their mask identities; 2) fully decomposing the 3D scene into discrete groups, e.g. representing different object instances to enable editing; 3) allowing for fast training and rendering, while not diluting the original 3D reconstruction quality.
Gaussian Grouping effectively leverages the dense 2D mask proposals of SAM, and lifts them to segment anything in the 3D scene via radiance fields rendering. In Sec 3.1, we first briefly review the 3D Gaussian Splatting method on radiance field rendering. We then detail the input data pre-processing steps and further describe the proposed Gaussian Grouping in Section 3.2. With the constructed 3D representation, finally, we show its advantages in downstream scene editing tasks by the efficient Local Gaussian Editing in Section 3.3.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting, as introduced by [15], encapsulates 3D scene information using a suite of 3D colored Gaussians. This technique has established its effectiveness in the reconstruction tasks, exhibiting high inference speeds and remarkable quality of reconstruction within timeframes on par with those of NeRF. Yet, its potential in scene understanding has not been thoroughly investigated. Our research reveals that 3D Gaussians hold considerable promise for the open-world and complex 3D scene understanding.
To represent the scene, each Gaussian’s property is characterized by a centroid $\mathbf{p}=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D size $\mathbf{s}\in\mathbb{R}^{3}$ in standard deviations, and a rotational quaternion $\mathbf{q}\in\mathbb{R}^{4}$. To allow fast $\alpha$-blending for rendering, an opacity value $\alpha\in\mathbb{R}$ and a color vector $\mathbf{c}$ are represented in the three degrees of spherical harmonics (SH) coefficients. These adjustable parameters are collectively symbolized by $S_{\Theta_{i}}$, where $S_{\Theta_{i}}=\{\mathbf{p}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}\}$ represents the set of parameters for the $i$-th Gaussian. Gaussian Splatting projects these 3D Gaussians onto the 2D image plane and implement differentiable rendering for each pixel.
@@
In this section, we describe the design of our Gaussian Grouping. To enable 3D Gaussians for fine-grained scene understanding, our key insight is that we preserve all attributes of the Gaussians (such as their position, color, opacity, and size) in their original setting, but add new Identity Encoding parameters (similar to the format of color modeling). This allows each Gaussian to be assigned to its represented instances or stuff in the 3D scene.
Figure 2: The method pipeline of our Gaussian Grouping contains three main steps: (a) We first prepare the input by deploying SAM to automatically generate masks in everything mode for each view independently. (b) Then, to obtain the consistent mask IDs across training views, we take a universal temporal propagation model [7] to associate the mask labels and generate a coherent multi-view segmentation. (c) With the prepared training input, we jointly learn all properties of the 3D Gaussians, including their group Identity Encoding, by differentiable rendering. Our encoding is supervised by the 2D Identity Loss, leveraging the coherent segmentation views, and a 3D Regularization loss. We use color the denote object IDs across frames for input views. We omit the rendering process for other Gaussian parameters and the density control part for simplicity, as it is inherited from [15].
Algorithm 1 Gaussian Grouping
$p\leftarrow$ SfM Points $\triangleright$ 3D Positions
$i\leftarrow 0$triangleright Iteration Count
while not converged do
    $L_{\text{id}}\leftarrow\lambda_{2d}\mathcal{L}_{\text{2d}}(\textbf{E}_{\text{id}},\hat{M})+\lambda_{3d}\mathcal{L}_{\text{3d}}(\textbf{e})$triangleright Identity Grouping Loss, Eq. 3
    if IsRefinementIteration($i$then
[…]
             RemoveGaussian()
          end if
          if $\nabla_{p}L>\tau_{p}$ then $\triangleright$ Densification
             if $\|S\|>\tau_{S}$ then $\triangleright$ Over-reconstruction
                SplitGaussian($p_{j},s_{j},\alpha_{j},c_{j},e_{j}$)
             $\triangleright$ Under-reconstruction
                CloneGaussian($p_{j},s_{j},\alpha_{j},c_{j},e_{j}$)
             end if
          end if
       end for
    end if
    $i\leftarrow i+1$
end while
The process of Gaussian Grouping is illustrated in Figure 2. We outline the pseudocode for our Gaussian Grouping in Algorithm 1, where we highlight the introduced core components in both red and bold texts.
3.2.1 (a) 2D Image and Mask Input
@@
To prepare the input for Gaussian Grouping, in Figure 2(a), we first deploy SAM to automatically generate masks for each image of the multi-view collection. The 2D masks are individually produced per image. Then, to assign each 2D mask a unique ID in the 3D scene, we need to associate the masks of the same identity across different views and obtain the total number $K$ of instances / stuff in the 3D scene.
3.2.2 (b) Identity Consistency across Views
@@
Instead of resorting to the cost-based linear assignment [44] during training, we treat the multi-view images of a 3D scene as a video sequence with gradually changing views. To achieve the 2D mask consistency across views, we employ a well-trained zero-shot tracker [7] to propagate and associate masks. This also provides the total number of mask identities in the 3D scene. We visualize associated 2D mask labels in Figure 2(b). Compared to the cost-based linear assignment proposed in [44], we find it simplifies the training difficulty while avoiding repeatedly computing the matching relation in each rendering iteration, resulting in over 60$\times$ speeding up. It also achieves better performance than the cost-based linear assignment, especially under the dense and overlapping masks by SAM. Besides, we show the robustness of our 3D masks association in Figure 5, where the 2D associated masks [7] from video contain obvious errors.
3.2.3 (c) 3D Gaussian Rendering and Grouping
@@
To generate 3D-consistent mask identities across views of the scene, we propose to group 3D Gaussians belonging to the same instance / stuff. In addition to the existing Gaussian properties, we introduce a new parameter, i.e., Identity Encoding, to each Gaussian. The Identity Encoding is a learnable and compact vector of length 16, which we find is sufficient to distinguish the different objects / parts in the scene with computation efficiency. During training, similar to Spherical Harmonic (SH) coefficients representing the color of each Gaussian, we optimize the introduced Identity Encoding vector to represent its instance ID of the scene. Note that different from the view-dependent appearance modeling of the scene, the instance ID is consistent across various rendering views. Thus, we set the SH degree of the Identity Encoding to 0, to only model its direct-current component. Unlike NeRF-based methods [44, 19, 16] designing extra semantic MLP layers, the Identity Encoding serves as the learnable property for each Gaussian to group the 3D scene.
@@
To optimize the introduced Identity Encoding of each Gaussian, in Figure 2(c), we render these encoded identity vectors into 2D images in a differentiable manner. We take the differentiable 3D Gaussian renderer from [15] and treat the rendering process similar to the color (SH coefficients) optimization in [15].
@@
3D Gaussian splatting adopts neural point-based $\alpha^{\prime}$-rendering [21, 20], where the influence weight $\alpha^{\prime}$ of each Gaussian can be evaluated in 2D for each pixel. Following [15], the influence of all Gaussians on a single pixel location is computed by sorting the Gaussians in depth order and blending $\mathcal{N}$ ordered points overlapping the pixels [30, 33]:
@@
3.2.4 (d) Grouping Loss
After associating 2D instance labels across each training view, suppose there are $K$ masks in total in the 3D scene. To group each 3D Gaussian by the instance /stuff mask identities, we design the grouping loss $\mathcal{L}_{\text{id}}$ for updating the Identity Encoding of Gaussians with two components:
@@
2. 3D Regularization Loss: To further boost the grouping accuracy of Gaussians, besides the standard cross-entropy loss for indirect 2D supervision, we also introduce an unsupervised 3D Regularization Loss to directly regularize the learning of Identity Encoding $e_{i}$. 3D Regularization Loss leverages the 3D spatial consistency, which enforces the Identity Encodings of the top $k$-nearest 3D Gaussians to be close in their feature distance. This allows the 3D Gaussians inside the 3D object, or heavily occluded (not visible in nearly all training views) during the point-based rendering (Eq. 1) to be supervised more sufficiently. In Eq. 3, we denote $F$ as softmax operation combined after linear layer $f$ (shared in computing the 2D Identity loss). We formalize the KL divergence loss with $m$ sampling points as,